OpenStudy (anonymous):

Factor completely: 2x2 - 2x - 12.

OpenStudy (anonymous):

28..

OpenStudy (anonymous):

factor...

OpenStudy (anonymous):

First I'd get rid of the 2: 2(x2 - x - 6) Then factor the rest: 2(x-3)(x+2)

OpenStudy (anonymous):

2(x - 3)(x + 2) x= 3 x= -2

OpenStudy (anonymous):

^^^^^^^^^^^ both of them are correct

OpenStudy (anonymous):

its telling me the correct answer is (2x+4) (x-3)

OpenStudy (anonymous):

That's weird, you are suppose to factor first.. I would talk to your teacher about that

OpenStudy (anonymous):

if u will take common from the first one (2x+4) (x-3) = 2(x+2)(x-3) both mean the same

OpenStudy (anonymous):

wow lol i didnt thnk of that sheg thanks